Magnesium hydroxide solid is added into 50 cm^3 of 1.5 mol dm^-3 nitric acid. The mixture is stirred until there is no more change observed. Calculate the mass of magnesium hydroxide needed to react completely with the nitric acid


1
Expert's answer
2021-08-19T18:27:04-0400

Mg(OH)2 + 2HNO3 = Mg(NO3)2 + 2H2O

According to the equation, n (HNO3) = 2 x n (Mg(OH)2)

Cn = n / V

n (Mg(OH)2) = 1.5 x 0.05 = 0.075 mol

n = m / M

m = n x M

M (Mg(OH)2) = 58.3 g/mol

m (Mg(OH)2) = 0.075 x 58.3 = 4.4 g